# Generalized Entropy Calibration for Inference with Partially Observed Data: A Unified Framework

### v1

@@ Abstract
Missing data is an universal problem in statistics. We develop a unified framework for estimating parameters defined by general estimating equations under a missing-at-random (MAR) mechanism, based on generalized entropy calibration weighting. We construct weights by minimizing a convex entropy subject to (i) balancing constraints on a data-adaptive calibration function, estimated using flexible machine-learning predictors with cross-fitting, and (ii) a debiasing constraint involving the fitted propensity score (PS) model. The resulting estimator is doubly robust, remaining consistent if either the outcome regression (OR) or the PS model is correctly specified, and attains the semiparametric efficiency bound when both models are correctly specified. Our formulation encompasses classical inverse probability weighting (IPW) and augmented IPW (AIPW) as special cases and accommodates a broad class of entropy functions. We illustrate the versatility of the approach in three important settings: semi-supervised learning with unlabeled outcomes, regression analysis with missing covariates, and causal effect estimation in observational studies. Extensive simulation studies and real-data applications demonstrate that the proposed estimators achieve greater efficiency and numerical stability than existing methods. In particular, the proposed estimator outperforms the classical AIPW estimator under the OR model misspecification.
Keywords: Calibration estimation, doubly robust, information projection, selection bias.
@@ 1 Introduction
Missing and incomplete data arise routinely in medical research, the social sciences, economics, and many other empirical disciplines. When the observed units are not representative of the target population, naïve complete-case analysis can produce biased and inefficient inference. A central goal of the missing-data literature is therefore to develop consistent and efficient estimators under plausible assumptions about the missingness mechanism (Little and Rubin, 2019; Kim and Shao, 2021).
Many modern data-analytic tasks can be viewed through the lens of partially observed data. In semi-supervised learning, the challenge is to exploit a large pool of unlabeled covariates for efficiency while guarding against covariate-shift bias (Chapelle et al., 2010; Zhang et al., 2019). In causal inference, systematic covariate imbalance between treatment groups must be corrected through weighting or matching (Hainmueller, 2012; Li et al., 2018). In regression with missing covariates, complete-case analysis is generally biased under missing at random (MAR; Ibrahim et al., 2005; Han and Wang, 2013). All three settings share a common structure: a parameter defined by an estimating equation that cannot be solved directly because part of the data vector is unobserved.
Two broad strategies exist. Outcome regression (OR) imputes the missing components via a conditional expectation model. Propensity-score (PS) weighting constructs inverse-probability weights (IPW). Doubly robust (DR) methods such as augmented IPW (AIPW; Robins et al., 1994) combine both and yield consistent estimators if either model is correctly specified. However, as demonstrated by Kang et al. (2007), standard DR estimators can be highly unstable with extreme weights. Targeted minimum loss-based estimation (TMLE; van der Laan and Rose, 2011) provides an alternative DR approach via a targeting step, but relies on a substitution estimator and does not directly produce calibration weights for the observed units. In parallel, entropy-based calibration methods choose weights by minimizing a convex divergence subject to balancing constraints (Zubizarreta, 2015; Chan et al., 2016; Zhao, 2019; Tan, 2020), achieving improved stability. These are closely related to covariate balancing methods in causal inference (Fan et al., 2022; Wang and Zubizarreta, 2020; Chattopadhyay et al., 2020; Hirshberg and Wager, 2021; Ben-Michael et al., 2021). However, existing calibration formulations do not directly yield a general doubly robust framework for arbitrary estimating equations with partially observed data.
In this paper, we develop a unified framework for analyzing partially observed data under MAR by combining generalized entropy calibration with doubly robust estimation and modern machine learning. The calibration weights are obtained by minimizing a generalized entropy subject to two sets of constraints: a balancing constraint based on a calibration function that approximates the optimal augmentation term, and a debiasing constraint incorporating a working PS model. To flexibly approximate the optimal calibration function, we embed cross-fitted machine-learning predictions into the calibration step, following the prediction-powered (Angelopoulos et al., 2023) and double machine learning (Chernozhukov et al., 2018) literatures.
The present work builds on the generalized entropy calibration framework of Kwon et al. (2025), which was developed for estimating finite population totals under design-based inference with known auxiliary variables. We extend their framework in three key directions. First, the target of inference is generalized from population totals to solutions of estimating equations $\mathbb{E}\{U(\theta;Z)\}=0$ under arbitrary MAR structures, which introduces a $\theta$-dependent calibration function requiring profile optimization (Algorithm 1) with no counterpart in the survey setting. Second, we establish double robustness and show that the proposed estimator achieves a strictly smaller asymptotic variance than AIPW under outcome model misspecification (Corollary 3), results that have no analog in the design-based framework where consistency follows from the known sampling mechanism. This variance-reduction property distinguishes our approach from both TMLE and the regularized calibrated estimation of Tan (2020), neither of which provides a similar mechanism under outcome-model misspecification. Third, we develop a cross-fitting procedure that integrates machine-learning predictions into the calibration constraints, and demonstrate the unifying scope of the approach across causal inference, semi-supervised learning, and missing covariate problems.
The main contributions are as follows. First, we formulate semi-supervised learning, regression with missing covariates, and causal inference within a single calibration-weighting framework for general estimating equations under MAR. Second, we developed a doubly robust generalized entropy calibration estimator that is doubly robust and locally efficient. Third, we developed a prediction-powered calibration. Cross-fitted machine-learning predictions enable flexible nonparametric calibration while maintaining valid large-sample theory. Simulations and real-data applications demonstrate substantial efficiency gains and improved stability, particularly under OR model misspecification.
@@
The remainder of the paper is organized as follows. Section 2 introduces the setup and reviews IPW and AIPW. Section 3 presents the proposed estimators. Section 4 establishes large-sample properties. Section 5 discusses computation. Section 6 illustrates applications to causal inference, semi-supervised learning, and missing covariates. Section 7 reports simulations, and Section 8 presents real-data applications. Some concluding remarks are made in Section 9. All the technical proofs are relegated to the supplementary material (SM).
2 Preliminaries and Existing Methods
Let $Z=(X^{\top},Y)^{\top}$ be a $(p+1)$-dimensional random vector where $X\in\mathbb{R}^{p}$ denotes covariates and $Y$ is the outcome variable. Suppose we observe independent and identically distributed copies $Z_{1},\ldots,Z_{N}$ of $Z$. For each unit $i$, decompose the full data vector as
$Z_{i}=\bigl(O_{i}^{\top},M_{i}^{\top}\bigr)^{\top},$ (2.1)
where $O_{i}$ is the always-observed subvector and $M_{i}$ is the potentially missing subvector. This decomposition encompasses several familiar settings:
  1. 1.
    Missing outcomes: $O_{i}=X_{i}$, $M_{i}=Y_{i}$;
  2. 2.
    Missing covariates: $O_{i}=Y_{i}$, $M_{i}=X_{i}$;
  3. 3.
    Partially missing covariates: $O_{i}=(X_{1i},Y_{i})^{\top}$, $M_{i}=X_{2i}$, where $X_{i}=(X_{1i}^{\top},X_{2i}^{\top})^{\top}$.
Each setting is developed in detail in Section 6.
Let $\delta_{i}=1$ if $M_{i}$ is observed and $\delta_{i}=0$ otherwise, and assume a missing-at-random (MAR) mechanism (Rubin, 1976): $\delta_{i}\perp M_{i}\mid O_{i}$. Let $\theta\in\mathbb{R}^{q}$ denote the parameter of interest, defined as the unique solution to the population estimating equation $\mathbb{E}\{U(\theta;Z)\}=0$, where $U(\theta;Z)$ is a $q\times 1$ estimating function. Without missing data, $\theta$ can be estimated by solving
$\hat{U}_{N}(\theta)\equiv\sum_{i=1}^{N}U(\theta;z_{i})=0.$ (2.2)
When $M_{i}$ is missing for some units, $U(\theta;z_{i})$ is not fully computable and (2.2) cannot be solved directly.
Under MAR, a common strategy is to model the response mechanism via a propensity-score (PS) model,
$P(\delta_{i}=1\mid O_{i})=\pi(O_{i};\phi),$ (2.3)
where $\phi$ is a finite-dimensional parameter estimated by maximum likelihood. The inverse-probability-weighted (IPW) estimator of $\theta$ solves
$\sum_{i=1}^{N}\frac{\delta_{i}}{\pi(O_{i};\hat{\phi})}\,U(\theta;z_{i})=0.$ (2.4)
@@ 2 Preliminaries and Existing Methods
To improve efficiency, Robins et al. (1994) proposed the augmented IPW (AIPW) estimator, which solves
$\sum_{i=1}^{N}\left[\frac{\delta_{i}\,U(\theta;z_{i})}{\pi(O_{i};\hat{\phi})}-\frac{\delta_{i}-\pi(O_{i};\hat{\phi})}{\pi(O_{i};\hat{\phi})}\,b(O_{i})\right]=0,$ (2.5)
where $b(O_{i})$ is an arbitrary function of $O_{i}$.
The choice of $b(O_{i})$ is of central importance. A particularly important target is
$b^{*}(\theta;O_{i})=\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\},$ (2.6)
the conditional expectation of the estimating function given $O_{i}$. If an OR model is specified for $\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\}$ and correctly estimated, then using (2.6) in the AIPW estimator yields an estimator that is approximately unbiased even when the PS model is misspecified. Consequently, the AIPW estimator enjoys a doubly robust property if either the PS model or the OR model is correctly specified, and locally efficient if both are correct (Robins et al., 1994).
However, the AIPW estimator relies on an explicit augmentation term $b(O_{i})$ and can be sensitive to extreme inverse-probability weights, motivating the calibration-based approach developed in Section 3.
3 Proposed Estimators
We propose an alternative to AIPW that constructs data-adaptive weights $\omega_{i}=\omega(O_{i})$ through a generalized entropy calibration procedure. Rather than using an explicit augmentation term, the method incorporates auxiliary information implicitly via calibration constraints, preserving double robustness while improving numerical stability.
We define the estimator $\hat{\theta}_{\omega}$ as the solution to the weighted estimating equation
$\sum_{i=1}^{N}\delta_{i}\,\hat{\omega}_{i}(\theta)\,U(\theta;z_{i})=0,$ (3.1)
where $\hat{\omega}_{i}(\theta)=\hat{\omega}(\theta;O_{i})$ are calibration weights that may depend on $\theta$, since the relevant auxiliary information for estimating $\theta$ may itself depend on $\theta$.
$\min_{\omega_{1},\ldots,\omega_{N}}\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i}),$ (3.2)
subject to the balancing constraint
$\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,b(O_{i})=\sum_{i=1}^{N}b(O_{i}),$ (3.3)
and the debiasing constraint
$\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,g(\hat{\pi}_{i}^{-1})=\sum_{i=1}^{N}g(\hat{\pi}_{i}^{-1}),$ (3.4)
[…]
Intuitively, if we use $b^{*}(O)=\mathbb{E}\left\{U(\theta;Z)\mid O\right\}$ in the balancing constraint and the OR model used to construct $b^{*}(O)$ is correct, balancing removes the leading bias term even when the PS model is wrong; conversely, if the PS model is correct, the debiasing constraint ensures consistency even when the OR model is wrong. These properties are formalized in Section 4.

### v2

# A Calibration Framework for Inference with Partially Observed Data

@@ Abstract
Many problems that are often studied separately, including causal inference, semi-supervised learning, and regression with incomplete covariates, share a common inferential structure: the parameter of interest is defined through an estimating equation involving a full-data vector, while part of that vector is unobserved. We study such problems under missing at random through generalized entropy calibration.
Our central claim is that calibration can be understood not only as a device for balancing covariates or stabilizing weights, but also as a general representation of augmentation. The proposed method constructs weights by minimizing a convex entropy subject to a balancing constraint, which incorporates outcome-related information through a calibration function, and a debiasing constraint, which is induced by a working propensity-score model.
This perspective yields a doubly robust estimator that attains the semiparametric efficiency bound when both working models are correct and, under correct propensity-score specification, can be more efficient than classical augmented inverse-probability weighting when the outcome model is misspecified. The framework also accommodates cross-fitting and applies across several partially observed-data problems within a common estimating-equation formulation. Simulations and real-data analyses illustrate the practical consequences of the theory.
Causal inference,
keywords:
[…]
@@ 1 Introduction
Many problems that are usually studied in separate literatures can be viewed as instances of a common inferential problem. In causal inference, one of the potential outcomes is unobserved for each unit. In semi-supervised learning, responses are observed only for a labeled subset. In regression with incomplete covariates, part of the predictor vector is missing. Although these settings differ in scientific interpretation, they share the same statistical structure: the parameter of interest is defined through a full-data estimating equation, but part of the full-data vector is unobserved. From this viewpoint, they are all problems of inference with partially observed estimating equations under missing at random (MAR); see, for example, Rubin (1976), Robins et al. (1994), Tsiatis (2006), Little and Rubin (2019), and Kim and Shao (2021).
The methodological responses to this problem have developed in different languages. In missing-data inference, the dominant themes have been outcome regression, inverse-probability weighting, and doubly robust augmentation (Robins et al., 1994; Rotnitzky and Robins, 1995; Scharfstein et al., 1999; Bang and Robins, 2005; Cao et al., 2009; Tsiatis, 2006). In survey sampling and nonresponse adjustment, calibration weighting provides a different formulation (Deville and Särndal, 1992; Deville, 2000; Wu and Sitter, 2001). In causal inference, balancing weights have become a central tool for covariate adjustment (Hainmueller, 2012; Imai and Ratkovic, 2014; Li et al., 2018; Zhao, 2019; Tan, 2020; Fan et al., 2022; Hirshberg and Wager, 2021; Ben-Michael et al., 2021). Related work has also emphasized the close connection between balancing and modeling perspectives on weighting (Chattopadhyay et al., 2020). Semi-supervised learning has developed its own efficiency-based perspective on the use of unlabeled data (Chakrabortty and Cai, 2018; Gronsbell and Cai, 2018; Zhang et al., 2019; Azriel et al., 2022). What remains less clear is how these strands fit together within a single estimating-equation framework.
A standard route through this problem starts from two familiar ideas. Outcome-regression (OR) methods replace the missing part of the data by a conditional mean or prediction, whereas propensity-score (PS) weighting corrects selective observation by reweighting the complete cases using inverse selection probabilities (Horvitz and Thompson, 1952). Doubly robust (DR) procedures combine these two strategies and remain consistent when either the OR model or the PS model is correctly specified (Robins et al., 1994; Rotnitzky and Robins, 1995; Scharfstein et al., 1999; Bang and Robins, 2005; Cao et al., 2009). This logic is usually written in additive form: one starts from an inverse-probability-weighted estimating equation and then adds an augmentation term. That representation is natural, but it is not the only one.
Our central claim is that augmentation can be represented implicitly through calibration. Rather than adding an explicit augmentation term, we construct weights by minimizing a generalized entropy subject to two constraints: a balancing constraint that incorporates outcome-related information through a calibration function, and a debiasing constraint induced by a working PS model. In this formulation, calibration is not merely a tool for moment balancing or weight stabilization. It becomes a general representation of augmentation itself. This places weighting, balancing, and doubly robust correction within a single framework for inference with partially observed estimating equations.
The present work builds on the generalized entropy calibration framework of Kwon et al. (2025), which was developed for estimating finite-population totals under design-based inference with known auxiliary variables. Our setting is different in both target and interpretation. Here the parameter is defined through a general estimating equation under MAR, the calibration function may depend on the target parameter itself, and estimation proceeds through profile optimization. These differences allow calibration to recover outcome-related information from the full sample while preserving the debiasing role of inverse-probability weighting. Recent work by Bruns-Smith et al. (2025) also highlights links between balancing weights and regression adjustment in linear settings; the present paper develops a broader calibration-based perspective for general estimating equations with partially observed data.
This perspective clarifies the geometry behind double robustness and efficiency. When the calibration function captures the relevant conditional mean structure, the balancing constraint plays the role usually assigned to augmentation. When the PS model is correctly specified, the debiasing constraint recovers the inverse-probability-weighting pathway. Double robustness then emerges from the interaction of these two constraint systems. The same perspective also clarifies why efficiency can improve relative to classical augmented inverse-probability weighting (AIPW). Standard AIPW augments through a working OR term, whereas the generalized entropy calibration estimator projects onto a larger space generated jointly by outcome-related and debiasing covariates. As a result, when the PS model is correct but the working OR model is misspecified, the proposed estimator can have strictly smaller asymptotic variance than classical AIPW, while under joint model correctness it attains the same semiparametric efficiency bound identified by Robins et al. (1994).
The same framework applies naturally across problems that are often treated as distinct. In causal inference, the partially observed component is a counterfactual outcome. In semi-supervised learning, it is an unlabeled response. In regression with incomplete covariates, it is a missing part of the predictor vector. The scientific meanings differ, but the inferential template is the same: one observes an always-observed component, a selectively observed component, and an estimating function defined on the full data. The examples in this paper are included not simply to demonstrate the reach of a new estimator, but to show that the same calibration logic recurs across several domains of modern statistics.
The framework also connects naturally to modern prediction-based inference. Because the balancing constraint requires calibration-function values for all units, including those with unobserved components, it is natural to estimate these quantities using cross-fitted predictions. This links the present approach to double machine learning (Chernozhukov et al., 2018) and the recent prediction-powered perspective of Angelopoulos et al. (2023). Here, however, machine learning plays a supporting role: it enriches the calibration space without replacing the underlying estimating-equation structure.
This paper makes three contributions. First, it formulates causal inference, semi-supervised learning, and regression with incomplete covariates within a common framework of inference with partially observed estimating equations under MAR. Second, it shows that calibration can be understood not only as a weighting device but as a general representation of augmentation, thereby linking calibration weighting to the logic of doubly robust inference. Third, it provides a geometric explanation of why this representation yields double robustness, recovers local efficiency under joint model correctness, and can improve upon classical AIPW when the PS model is correct but the working OR model is misspecified.
@@
The remainder of the paper is organized as follows. Section 2 introduces the general setup and reviews OR, IPW, and AIPW. Section 3 presents generalized entropy calibration as a representation of augmentation. Section 4 develops the main asymptotic implications of this viewpoint. Section 5 discusses computation and cross-fitting. Section 6 shows how the framework specializes to causal inference, semi-supervised learning, and regression with missing covariates. Section 7 reports simulation results, and Section 8 presents real-data applications. Some concluding remarks are given in Section 9. All technical proofs are deferred to the supplementary material.
2 Basic Setup
We consider a general partially observed-data setting in which the full data vector for unit $i$ is $Z_{i}=(O_{i}^{\top},M_{i}^{\top})^{\top}$, where $O_{i}$ denotes the subvector of $Z_{i}$ that is always observed and $M_{i}$ denotes the subvector subject to missingness. Let $\delta_{i}\in\{0,1\}$ indicate whether $M_{i}$ is observed, so that the observed data are $(O_{i},\delta_{i},\delta_{i}M_{i})$. The parameter of interest $\theta\in\mathbb{R}^{p}$ is defined through the full-data estimating equation
$\mathbb{E}\{U(\theta;Z_{i})\}=0,$
where $U(\theta;Z_{i})$ is a known estimating function. This setup encompasses a wide range of problems, including causal inference, semi-supervised learning, and regression with incomplete covariates. The statistical task is to recover the full-data estimating equation, and hence $\theta$, from partially observed data under assumptions on the missingness mechanism.
To identify $\theta$ from the observed data, we assume that the missingness mechanism is missing at random (MAR), that is,
$\delta_{i}\perp M_{i}\mid O_{i}.$
Thus, conditional on the always-observed subvector $O_{i}$, the indicator of whether $M_{i}$ is observed does not depend further on the missing subvector itself. Let
$\pi(O_{i})=P(\delta_{i}=1\mid O_{i})$
denote the propensity score, and assume positivity, $0<\pi(O_{i})\leq 1$ almost surely. Under these conditions, the full-data estimating equation can be recovered from the observed data through inverse-probability weighting, since
$\mathbb{E}\!\left\{\frac{\delta_{i}}{\pi(O_{i})}\,U(\theta;Z_{i})\right\}=\mathbb{E}\{U(\theta;Z_{i})\}.$
This identity underlies the IPW estimator and also serves as a starting point for augmented and calibration-based constructions.
Under MAR, a natural approach is inverse-probability weighting (IPW), which reconstructs the full-data estimating equation by reweighting the complete observations. Specifically,
$\mathbb{E}\!\left\{\frac{\delta_{i}}{\pi(O_{i})}\,U(\theta;Z_{i})\right\}=\mathbb{E}\{U(\theta;Z_{i})\}$
suggests estimating $\theta$ by solving
$\frac{1}{N}\sum_{i=1}^{N}\frac{\delta_{i}}{\pi(O_{i};\hat{\phi})}\,U(\theta;Z_{i})=0,$
where $\pi(O_{i};\phi)$ is a working propensity-score model and $\hat{\phi}$ is an estimator of $\phi$. If the propensity-score model is correctly specified, the resulting IPW estimator is consistent for $\theta_{0}$. However, because IPW relies only on the response mechanism, it can be inefficient and numerically unstable when some estimated propensity scores are small. This limitation motivates augmented constructions that incorporate outcome-related information in addition to inverse-probability weighting.
A standard refinement of IPW is augmented inverse-probability weighting (AIPW), which combines weighting with an explicit outcome-regression adjustment. Let $b(\theta;O_{i})$ denote a working model for
$b^{*}(\theta;O_{i})=\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\},$
the conditional mean of the full-data estimating function given the always-observed subvector $O_{i}$. The AIPW estimator is obtained by solving
$\frac{1}{N}\sum_{i=1}^{N}\left[b(\theta;O_{i})+\frac{\delta_{i}}{\pi(O_{i};\hat{\phi})}\bigl\{U(\theta;Z_{i})-b(\theta;O_{i})\bigr\}\right]=0.$
This construction is doubly robust: it remains consistent for $\theta_{0}$ if either the propensity-score model or the outcome-regression model is correctly specified. In addition, when both working models are correct and $b(\theta;O_{i})=b^{*}(\theta;O_{i})$, the estimator attains the semiparametric efficiency bound. Thus, AIPW improves upon IPW by incorporating outcome-related information through an explicit augmentation term (Robins et al., 1994).
However, the AIPW estimator relies on an explicit augmentation term $b(\theta;O_{i})$ and can be sensitive to extreme inverse-probability weights, motivating the calibration-based approach developed in Section 3.
3 Proposed Estimators
The augmented inverse-probability weighted estimator improves upon IPW by adding an explicit augmentation term $b(O_{i})$ to the estimating equation. Our starting point is that this augmentation step can instead be represented implicitly through calibration. Rather than inserting $b(O_{i})$ directly into the estimating equation, we construct weights so that the weighted complete cases reproduce both outcome-related information and response-mechanism information. This leads to a calibration-based formulation of doubly robust estimation in which weighting, augmentation, and balance are integrated within a single framework.
More specifically, the proposed weights are designed to satisfy two goals. The first is to balance a calibration function
$b(O)=\bigl(1,\tilde{b}(O)^{\top}\bigr)^{\top}$
that approximates the conditional mean
$b^{*}(\theta;O)=E\{U(\theta;Z)\mid O\},$
thereby capturing the same information that appears in the classical augmentation term. We assume throughout that the first component of $b(O)$ is an intercept term. Consequently, the balancing constraint enforces not only moment balance for the nonconstant components of $b(O)$, but also total-weight calibration. The second goal is to incorporate a working propensity-score model through an additional debiasing constraint, so that the resulting weights recover the inverse-probability weighting structure when the PS model is correctly specified. In this sense, the proposed construction provides a calibration representation of the usual OR/IPW synthesis.
We therefore define the estimator $\hat{\theta}_{\omega}$ as the solution to the weighted estimating equation
$\sum_{i=1}^{N}\delta_{i}\,\hat{\omega}_{i}(\theta)\,U(\theta;z_{i})=0,$ (1)
where $\hat{\omega}_{i}(\theta)=\hat{\omega}(\theta;O_{i})$ denotes a calibration weight assigned to unit $i$. The dependence on $\theta$ is essential: because the target estimating function itself depends on $\theta$, the auxiliary information used to calibrate the weights may also depend on $\theta$.
$\min_{\omega_{1},\ldots,\omega_{N}}\sum_{i=1}^{N}\delta_{i}\,G(\omega_{i}),$ (2)
subject to two constraints, each corresponding to one side of the doubly robust construction.
The first is the balancing constraint
$\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,b(O_{i})=\sum_{i=1}^{N}b(O_{i}),$ (3)
which requires the weighted complete cases to reproduce the empirical moments of the calibration function $b(O)$. Because the first component of $b(O)$ is an intercept, (3) also implies the total-weight calibration identity
$\sum_{i=1}^{N}\delta_{i}\,\omega_{i}=N.$
When $b(O)$ approximates $b^{*}(\theta;O)=\mathbb{E}\{U(\theta;Z)\mid O\}$, this constraint imports outcome-related information from the full sample into the respondent sample and plays the same role as the augmentation term in AIPW.
The second is the debiasing constraint
$\sum_{i=1}^{N}\delta_{i}\,\omega_{i}\,g(\hat{\pi}_{i}^{-1})=\sum_{i=1}^{N}g(\hat{\pi}_{i}^{-1}),$ (4)
@@ 3 Proposed Estimators
$\omega_{i}^{*}=g^{-1}(\lambda^{\top}s_{i}),$
for an appropriate dual variable $\lambda$; see Section 3.1. Consequently, if calibration is imposed only through the debiasing constraint and the PS model is correct, then
$\omega_{i}^{*}=g^{-1}\{g(\pi_{i}^{-1})\}=\pi_{i}^{-1},$
which exactly recovers the usual IPW weights. Thus, the debiasing covariate $g(\hat{\pi}_{i}^{-1})$ is not an ad hoc choice; it is the form required for the calibration weights to inherit the IPW structure under a correct PS model.
The two constraints therefore play distinct but complementary roles. The balancing constraint injects outcome-related information through the calibration function $b(O)$, whereas the debiasing constraint incorporates the fitted propensity-score model and protects consistency under correct PS specification. Together they provide a calibration-based analogue of the classical doubly robust construction. In particular, if $b(O)$ equals $b^{*}(\theta;O)$, or is consistently estimated from a correct OR model, then balancing removes the leading bias term even when the PS model is misspecified. Conversely, if the PS model is correct, the debiasing constraint ensures that the calibration estimator inherits the consistency of IPW even when the OR model is misspecified. These properties are formalized in Section 4.